OpenStudy (anonymous):

Please help!!!!! MEDAL AND FAN!!!!! I think the answer is 75 degrees. Is it?

OpenStudy (anonymous):

This is the picture.

OpenStudy (anonymous):

20

OpenStudy (abb0t):

How did you get 75? Can you show us?

OpenStudy (anonymous):

not buying 75 degrees

OpenStudy (anonymous):

Well my first thought was that it was 115 but I got that wrong. 180-105=75 So I thought maybe that was right......

OpenStudy (anonymous):

do you know that the lines are diagonals? Honestly, they can't be. are the lines the same length?

OpenStudy (anonymous):

Then Idk how to solve this problem

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

Ohhh I think I get it...... maybe......

OpenStudy (anonymous):

130?

OpenStudy (anonymous):

we know that m<1 + m<1 + m<2 + m<2 = 360 m<2 = (125 +105)/2 = 115 so m<1 = ?

OpenStudy (anonymous):

125+105=230/2=115

OpenStudy (anonymous):

So the answer is 115

OpenStudy (anonymous):

go back and take a closer look...

OpenStudy (anonymous):

and remember what you want to solve

OpenStudy (freckles):

hint: straight lines have angle 180 degrees

OpenStudy (anonymous):

lol that's what m<2 is

OpenStudy (anonymous):

65?

OpenStudy (freckles):

:) bingo

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

Thanks so much.

OpenStudy (anonymous):

@ameenbriscoe13 y r u steel looking at dis question its done she opened a new question